OpenStudy (anonymous):

The length of a rectangle is 24 units. Can the perimeter x of the rectangle be 60 units when its width y is 11 units?

Mehek (mehek14):

let's check \(2*24+2*11=60\\2*24=48\\2*11=22\\48+22=70\)

Mehek (mehek14):

so it can't be 60